Find the number of ways 3 boys and 4 girls can be arranged in a row if there are no restrictions?
Multiple choice answers:
a. 5,040 Correct
b. 5,249
c. 2,789
d. 2,040

Answer by ikleyn(52776)

.

They can be arranged by (4+3)! = 7! = 7*6*5*4*3*2*1 = 5040 ways.

Solved.

-----------------

On Permutations,  see introductory lessons
    - Introduction to Permutations
    - PROOF of the formula on the number of Permutations
    - Simple and simplest problems on permutations
    - Special type permutations problems
    - Problems on Permutations with restrictions

    - OVERVIEW of lessons on Permutations and Combinations
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.